# Quantum measurement tomography with mini-batch stochastic gradient descent

### v1

@@ Abstract
Drawing inspiration from gradient-descent methods developed for data processing in quantum state tomography [Quantum Sci. Technol. 10 045055 (2025)] and quantum process tomography [Phys. Rev. Lett. 130, 150402 (2023)], we introduce stochastic gradient descent (SGD) algorithms for fast quantum measurement tomography (QMT), applicable to both discrete- and continuous-variable quantum systems—thus completing the tomography trio. A measurement device or detector in a quantum experiment is characterized by a set of positive operator-valued measure (POVM) elements; the goal of QMT is to estimate these operators from experimental data. To ensure physically valid (positive and complete) POVM reconstructions, we propose two distinct parameterization schemes within the SGD framework: one leveraging optimization on a Stiefel manifold and one based on Hermitian operator normalization via eigenvalue scaling. Within the SGD-QMT framework, we further investigate two loss functions: mean squared error, equivalent to L2 or Euclidean norm, and average negative log-likelihood, inspired by maximum likelihood estimation. We benchmark performance against state-of-the-art constrained convex optimization methods. Numerical simulations demonstrate that, compared to standard methods, our SGD-QMT algorithms offer significantly lower computational cost, superior reconstruction fidelity, and enhanced robustness to noise. We make a Python implementation of the SGD-QMT algorithms publicly available at github.com/agtomo/SGD-QMT.
@@ I Introduction
In recent years, the field of quantum information science has witnessed substantial progress in the development of quantum characterization, verification, and validation (QCVV) protocols [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. These protocols encompass a wide spectrum of techniques, generally organized into four major categories [1]: experimental device characterization, tomographic methods, randomized benchmarking, and task-specific benchmarking. Collectively, they provide essential tools for ensuring the reliable and scalable operation of quantum hardware. In particular, QCVV protocols not only serve as benchmarks for assessing device performance but also as diagnostics for identifying and quantifying error mechanisms, and as practical guides for optimizing control, calibration, and error-mitigation strategies, thereby laying the foundation for advancing from noisy intermediate-scale quantum (NISQ) systems toward scalable, fault-tolerant quantum computation [18, 19, 20, 21, 22]. The continued refinement of QCVV techniques is thus critical for advancing quantum technologies from proof-of-principle demonstrations toward robust, large-scale implementations.
@@
When applying QCVV techniques to quantum processors, the operation of the processor can, in general, be understood in terms of three fundamental components [23, 24, 25]: quantum states (inputs), quantum processes or gates (operations), and measurement readouts or detection (outputs); see Fig. 1. Each component in this trio plays a vital role in the functionality of the processor, and errors or imperfections in any part can lead to incorrect or unreliable computational outcomes. Therefore, it is extremely important to have efficient strategies for comprehensive characterization of the whole trio.
Among QCVV tools for comprehensive characterization of quantum systems, tomographic techniques are some of the most widely used. These techniques include: (i) quantum state tomography (QST) [26, 27, 28, 29, 30], which reconstructs the density matrix $\rho$ describing an unknown quantum state; (ii) quantum process tomography (QPT) [31, 32, 33, 34, 35, 36, 37], which estimates the action of a quantum process through a process matrix $\chi$; and (iii) quantum measurement tomography (QMT) [38, 39, 40, 41, 42, 43], also referred to as quantum detector tomography, which characterizes the measurement apparatus modeled by a set of positive operator-valued measure (POVM) elements $\{\Pi_{i}\}$. All these three tomographic techniques are statistical procedures that rely on experimental data and an estimator—an algorithm that maps the collected data to an estimate of the underlying state, process, or detector. In this article, we put forward and benchmark algorithms for such data processing in the case of QMT, inspired by earlier work on that topic for QST and QPT.
@@
$\text{QMT:}\hskip 2.84544pt\left\{\Pi_{i}\in\mathbb{C}^{d\times d}\right\}_{i=1}^{k}\hskip 5.69046pt\text{s.t.}\hskip 5.69046pt\sum_{i=1}^{k}\Pi_{i}=I\hskip 2.84544pt\&\hskip 2.84544pt\Pi_{i}\geq 0,$ (3)
@@ III.1.1 Full-rank random POVM sets
Figure 2: Time complexity and performance metrics of different QMT data-processing methods for random full-rank POVM sets. The upper and lower lines for each marker are error bars denoting one standard deviation. The first row shows runtime (in seconds, on a $\log_{10}$ scale), while the second and third rows report the Frobenius distance and the Wasserstein distance, respectively, for the reconstructed POVMs (both also on a $\log_{10}$ scale). Each metric on the $y$ axes is averaged over multiple randomly generated POVM sets [20 for ($N\leq 4$ & $K\leq 16$), 5–10 for ($N=5,6$ & $K=32$)] of the indicated size. The first column corresponds to QMT with $k=8$ POVM elements, while second and third columns correspond to $k=16$ and $k=32$, respectively. In all cases, the $x$ axis denotes the number of qubits, with the maximum number of iterations fixed at 3000.
Figure 2 summarizes the performance of the SGD-QMT methods HONEST-MSE (teal), HONEST-MLE (orange), SM-MSE (red), and SM-MLE (green), and compares them against the state-of-the-art CCO-CVX approach (black). In all figure panels, the $x$ axis denotes the number of qubits. The first row of Fig. 2 shows the total computational time required to finish 3000 iterations and reconstruct the full POVM set. The second and third rows report the corresponding Frobenius norm and Wasserstein distance, respectively, which serve as benchmarking metrics for assessing the reconstruction quality. All three metrics on the $y$ axes are averaged over multiple randomly generated POVM sets of the specified sizes for each column.
@@
However, this dramatic gain in computational efficiency does come at the expense of slightly worse reconstruction quality, as evidenced by the Frobenius and Wasserstein metrics (the second and third rows in Fig. 2), where SGD-QMT have slightly or clearly lower accuracy, in some cases approaching that of CCO-CVX. The gap between SGD-QMT and CCO-CVX is larger in terms of the average Wasserstein distance, but that distance still consistently falls within the range of $10^{-5}-10^{-3}$ or better for the SGD-QMT methods, ensuring sufficiently high accuracy. Note that the maximum number of iterations for all SGD-QMT methods in all cases presented in Fig. 2 is set to 3000 for consistent comparison. The estimation accuracy can be further improved by increasing the iteration count for high-dimensional systems. These results highlight the scalability advantage of SGD-QMT methods, particularly in regimes where traditional convex optimization becomes computationally prohibitive.
While the runtime of all SGD-QMT algorithms are nearly identical, it is clear from Fig. 2 that, in almost all scenarios, the HONEST parameterization yields substantially better reconstruction quality, with Frobenius norm and Wasserstein distance values smaller than those obtained with the SM parameterization. Furthermore, the MSE loss function yields better results than the MLE loss function for the HONEST parameterization for smaller qubit numbers. For larger qubit numbers, the two loss functions do not differ much in the results for the HONEST parameterization, though we note that the MSE loss function seems to suit the SM parameterization better than the MLE loss function across most qubit and POVM element numbers.
@@ III.1.2 Rank-1 Pauli projective measurements
Next, we evaluate the performance of our SGD-QMT algorithms for Pauli projective measurements, which are described by $2^{N}$ rank-1 POVM elements (projectors), computed as an outer product of normalized eigenvectors of given $N$-qubit Pauli matrices. The results shown in Fig. 3 summarize the comparison of the SGD-QMT methods—HONEST-MSE (cyan), HONEST-MLE (orange), SM-MSE (red), and SM-MLE (green)—with the CCO-CVX method (purple, up to four qubits) in terms of three performance metrics, as a function of the number of qubits: average computational time [Fig. 3(a)], average Frobenius norm error [Fig. 3(b)], and average Wasserstein distance [Fig. 3(c)], each obtained by repeating the reconstruction over multiple randomly chosen Pauli projective measurements. For consistency, the maximum number of iterations in all cases is fixed at 1000.
The results in Fig. 3(a) indicate that the computational time complexity across all SGD-QMT algorithms remains nearly identical (as also observed in the case of full-rank random POVM sets in Fig. 2). However, the corresponding reconstruction quality, as quantified by the Frobenius norm and Wasserstein distance in Figs. 3(b) and 3(c), exhibits substantial variation, with differences spanning several orders of magnitude. For two and three qubits, CCO-CVX outperforms SGD-QMT in terms of runtime, requiring an order of magnitude less time while also achieving higher reconstruction quality. However, the efficiency of CCO-CVX deteriorates rapidly beyond three qubits, where it requires orders of magnitude more time than SGD-QMT methods while providing comparable reconstruction accuracy. We report CCO-CVX results only up to four qubits, as for the five-qubit system CCO-CVX failed to produce a solution even after one hour. In contrast, for random POVM sets, CCO-CVX was able to obtain a solution in roughly 15 minutes, as shown in Fig. 2. This significantly limits the practical usability of CVX for systems larger than four qubits. Among the tested SGD-QMT methods, the HONEST-MLE algorithm consistently achieves the highest reconstruction accuracy across all qubit numbers, delivering superior performance without incurring additional computational cost or requiring more iterations. This is different than the results for full-rank random POVM sets in Fig. 2, where HONEST-MSE performed best in terms of reconstruction accuracy, although HONEST-MLE generally was almost just as good.
@@ III.2.2 Photon counting
Photon counting in CV systems is mathematically equivalent to projective measurements in the computational basis in multi-qubit DV systems. The key distinction, however, lies in the choice of probe states, which are fundamentally distinct in the two cases. Photon counting is modeled with $k=\dim(\mathcal{H})$ number of rank-1 POVM elements, each of dimension $k\times k$, defined as

### v2

@@ Abstract
Drawing inspiration from gradient-descent methods developed for data processing in quantum state tomography [Quantum Sci. Technol. 10, 045055 (2025)] and quantum process tomography [Phys. Rev. Lett. 130, 150402 (2023)], we introduce stochastic gradient descent (SGD) algorithms for fast quantum measurement tomography (QMT), applicable to both discrete- and continuous-variable quantum systems—thus completing the tomography trio. A measurement device or detector in a quantum experiment is characterized by a set of positive operator-valued measure (POVM) elements; the goal of QMT is to estimate these operators from experimental data. To ensure physically valid (positive and complete) POVM reconstructions, we propose two distinct parameterization schemes within the SGD framework: one leveraging optimization on a Stiefel manifold and one based on Hermitian operator normalization via eigenvalue scaling. Within the SGD-QMT framework, we further investigate two loss functions: mean squared error, equivalent to L2 or Euclidean norm, and average negative log-likelihood, inspired by maximum likelihood estimation. We benchmark performance against state-of-the-art constrained convex optimization methods. Numerical simulations demonstrate that, compared to standard methods, our SGD-QMT algorithms offer significantly lower computational cost, superior reconstruction fidelity, and enhanced robustness to noise. We make a Python implementation of the SGD-QMT algorithms publicly available at github.com/agtomo/SGD-QMT.
@@ I Introduction
In recent years, the field of quantum information science has witnessed substantial progress in the development of quantum characterization, verification, and validation (QCVV) protocols [31, 19, 58, 40, 64, 52, 63, 10, 32, 70, 74, 15, 51, 68, 48, 57, 33, 49]. These protocols encompass a wide spectrum of techniques, generally organized into four major categories [31]: experimental device characterization, tomographic methods, randomized benchmarking, and task-specific benchmarking. Collectively, they provide essential tools for ensuring the reliable and scalable operation of quantum hardware. In particular, QCVV protocols not only serve as benchmarks for assessing device performance but also as diagnostics for identifying and quantifying error mechanisms, and as practical guides for optimizing control, calibration, and error-mitigation strategies, thereby laying the foundation for advancing from noisy intermediate-scale quantum (NISQ) systems toward scalable, fault-tolerant quantum computation [56, 28, 20, 5, 75]. The continued refinement of QCVV techniques is thus critical for advancing quantum technologies from proof-of-principle demonstrations toward robust, large-scale implementations.
@@
When applying QCVV techniques to quantum processors, the operation of the processor can, in general, be understood in terms of three fundamental components [53, 17, 29, 21]: quantum states (inputs), quantum processes or gates (operations), and measurement readouts or detection (outputs); see Fig. 1. Each component in this trio plays a vital role in the functionality of the processor, and errors or imperfections in any part can lead to incorrect or unreliable computational outcomes. Therefore, it is extremely important to have efficient strategies for comprehensive characterization of the whole trio.
Among QCVV tools for comprehensive characterization of quantum systems, tomographic techniques are some of the most widely used. These techniques include: (i) quantum state tomography (QST) [27, 66, 30, 60, 4], which reconstructs the density matrix $\rho$ describing an unknown quantum state; (ii) quantum process tomography (QPT) [14, 54, 39, 61, 24, 25, 41], which estimates the action of a quantum process through a process matrix $\chi$; and (iii) quantum measurement tomography (QMT) [22, 47, 46, 11, 72, 6], also referred to as quantum detector tomography, which characterizes the measurement apparatus modeled by a set of positive operator-valued measure (POVM) elements $\{\Pi_{i}\}$. All these three tomographic techniques are statistical procedures that rely on experimental data and an estimator—an algorithm that maps the collected data to an estimate of the underlying state, process, or detector. In this article, we put forward and benchmark algorithms for such data processing in the case of QMT, inspired by earlier work on that topic for QST [26] and QPT [3].
@@
$\text{QMT:}\hskip 2.84544pt\left\{\Pi_{i}\in\mathbb{C}^{d\times d}\right\}_{i=1}^{k}\hskip 2.84544pt\text{s.t.}\hskip 2.84544pt\sum_{i=1}^{k}\Pi_{i}=I\hskip 2.84544pt\&\hskip 2.84544pt\Pi_{i}\geq 0\hskip 1.42271pt\forall i,$ (3)
@@ II.5.3 Performance metrics
It is worth noting that one may also adopt the conventional readout fidelity metric—particularly relevant for measurements in the computational basis—which is commonly defined as the detector’s average success probability for correctly identifying computational basis states $|e_{i}\rangle$, with $|e_{i}\rangle\in\{|0\rangle,|1\rangle\}^{\otimes N}$. This metric is widely used for calibrating measurement hardware. However, we emphasize that readout fidelity only captures information about the diagonal entries of the reconstructed POVM elements and therefore offers a limited view of the measurement apparatus model. For this reason, we adopt the Frobenius (and Wasserstein) norm as our benchmarking metric, as it provides a global measure of discrepancy by accounting for contributions from all elements of the POVM matrices.
@@ III.1.1 Full-rank random POVM sets
Figure 2: Time complexity and performance metrics of different QMT data-processing methods for random full-rank POVM sets. The upper and lower lines for each marker are error bars denoting one standard deviation. The first row shows runtime in seconds, while the second and third rows report the Frobenius distance and the Wasserstein distance, respectively, for the reconstructed POVMs. Each metric on the $y$ axes is averaged 15 randomly generated POVM sets of the indicated size. The first column corresponds to QMT with $k=8$ POVM elements, while second and third columns correspond to $k=16$ and $k=32$, respectively. In all cases, the $x$ axis denotes the number of qubits, with the maximum number of iterations fixed at 2000. The data points for given qubit is confined within corresponding faint colored vertical band to avoid value overlap.
Figure 2 summarizes the performance of the SGD-QMT methods HONEST-MSE (teal), HONEST-MLE (orange), SM-MSE (red), and SM-MLE (green), and compares them against the state-of-the-art CCO-CVX approach (black). In all figure panels, the $x$-axis represents the number of qubits, with each value highlighted by a faint colored vertical band. All data points corresponding to a given qubit number are confined within the associated vertical band, providing clear visual separation and preventing overlap between values. The first row of Fig. 2 shows the total computational time (in seconds) required to finish 2000 iterations and reconstruct the full POVM set. The second and third rows report the corresponding Frobenius norm and Wasserstein distance, respectively, which serve as benchmarking metrics for assessing the reconstruction quality. All three metrics on the $y$ axes are averaged over $15$ randomly generated POVM sets of the specified sizes for each column.
@@
However, this dramatic gain in computational efficiency does come at the expense of slightly worse reconstruction quality, as evidenced by the Frobenius and Wasserstein metrics (the second and third rows in Fig. 2), where SGD-QMT have slightly or clearly lower accuracy, in some cases approaching that of CCO-CVX. The gap between SGD-QMT and CCO-CVX is larger in terms of the average Wasserstein distance, but that distance still consistently falls within the range of $10^{-5}-10^{-3}$ or better for the SGD-QMT methods, ensuring sufficiently high accuracy for most practical purposes. Note that the maximum number of iterations for all SGD-QMT methods in all cases presented in Fig. 2 is set to 2000 for consistent comparison. The estimation accuracy can be further improved by increasing the iteration count for high-dimensional systems. These results highlight the scalability advantage of SGD-QMT methods, particularly in regimes where traditional convex optimization becomes computationally prohibitive.
While the runtime of all SGD-QMT algorithms are nearly identical, it is clear from Fig. 2 that, in almost all scenarios, the HONEST parameterization yields substantially better reconstruction quality, with Frobenius norm and Wasserstein distance values smaller than those obtained with the SM parameterization. Furthermore, for the SM case, the MSE loss function outperforms the MLE loss function, whereas under the HONEST parameterization, the MSE and MLE loss functions exhibit nearly identical performance.
@@ III.1.2 Rank-1 Pauli projective measurements
Next, we evaluate the performance of our SGD-QMT algorithms for Pauli projective measurements, which are described by $2^{N}$ rank-1 POVM elements (projectors), computed as an outer product of normalized eigenvectors of given $N$-qubit Pauli matrices. The results shown in Fig. 3 summarize the comparison of the SGD-QMT methods—HONEST-MSE (cyan), HONEST-MLE (orange), SM-MSE (red), and SM-MLE (green)—with the CCO-CVX method (purple, up to four qubits) in terms of three performance metrics, as a function of the number of qubits: average computational time [Fig. 3(a)], average Frobenius norm error [Fig. 3(b)], and average Wasserstein distance [Fig. 3(c)], each obtained by repeating the reconstruction over multiple randomly chosen Pauli projective measurements. For consistency, the maximum number of iterations in all cases is fixed at 1000 and can be increased further for further improvement.
The results in Fig. 3(a) indicate that the computational time complexity across all SGD-QMT algorithms remains nearly identical (as also observed in the case of full-rank random POVM sets in Fig. 2). However, the corresponding reconstruction quality, as quantified by the Frobenius norm and Wasserstein distance in Figs. 3(b) and 3(c), exhibits substantial variation, with differences spanning several orders of magnitude. For two and three qubits, CCO-CVX outperforms SGD-QMT in terms of runtime, requiring an order of magnitude less time while also achieving higher reconstruction quality. However, the efficiency of CCO-CVX deteriorates rapidly beyond three qubits, where it requires orders of magnitude more time than SGD-QMT methods while providing comparable reconstruction accuracy with HONEST-MLE. We report CCO-CVX results only up to four qubits, as for the five-qubit system CCO-CVX failed to produce a solution even after one hour. In contrast, for random POVM sets, CCO-CVX was able to obtain a solution in roughly 15 minutes, as shown in Fig. 2. This significantly limits the practical usability of CVX for systems larger than four qubits. Among the tested SGD-QMT methods, the HONEST-MLE algorithm consistently achieves the highest reconstruction accuracy across all qubit numbers, delivering superior performance without incurring additional computational cost or requiring more iterations. This is different than the results for full-rank random POVM sets in Fig. 2, where HONEST-MSE performed best in terms of reconstruction accuracy, although HONEST-MLE generally was almost just as good.
@@ III.2.2 Photon counting
Photon counting in CV systems—also known as photon number resolving detectors, is mathematically equivalent to projective measurements in the computational basis in multi-qubit DV systems. The key distinction, however, lies in the choice of probe states, which are fundamentally distinct in the two cases. Photon counting is modeled with $k=\dim(\mathcal{H})$ number of rank-1 POVM elements, each of dimension $k\times k$, defined as